OpenStudy (anonymous):

Which of the following is the solution to the equation 25^(z + 2) = 125 ? z = 5.5 z = 3.5 z = –2.5 z = –0.5

OpenStudy (abb0t):

Take the natural log of the equation. HINT: \[\ a^n=ln (a)^n = n \ln(a)\] where 'a' is a constant.

OpenStudy (abb0t):

Then, use algebra to rearrange and solve for z.

OpenStudy (anonymous):

Could you work through the steps for me?

OpenStudy (anonymous):

please?

OpenStudy (abb0t):

I gave you a good amount of tip and hints. Try and work it out yourself first. If you are wrong, there are no penalties. I will help guide you in the right direction.

OpenStudy (anonymous):

z= 5.5 ?

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

?

OpenStudy (abb0t):

Not quite. How did you get that?

OpenStudy (anonymous):

well I got down to z= 11/2

OpenStudy (abb0t):

Your first step should be to take the natural log of everything and take the form: \[(z+2)\ln25 = \ln125\]

OpenStudy (abb0t):

Your next step should be to divide by...

OpenStudy (anonymous):

1n25 ? since its outside the parenthesis?

OpenStudy (abb0t):

Correct! Now you have: \[z+2 = \frac{ \ln125 }{ \ln25 }\]

OpenStudy (abb0t):

Now, what can you do next to solve for z?

OpenStudy (anonymous):

add 2 to each side

OpenStudy (abb0t):

Well, not add 2, but instead subtract. If you add 2 to both sides, you get z+4, and you're solving for z = #

OpenStudy (anonymous):

oh, what do we do from there?

OpenStudy (abb0t):

Subtract 2 from both sides and thats your answer. Plug it into a calculator to get your answer.

OpenStudy (anonymous):

z= -2.5 ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

hello can someone confirm this???????????????????????????????????

OpenStudy (anonymous):

\[25=5^2\] \[125=5^3\]

OpenStudy (anonymous):

\[25^{z+2}=5^{2z+4}=5^3\] \[2z+4=3\] \[2z=-1\] \[z=-\frac{1}{2}\]

OpenStudy (anonymous):

i have no idea what that log stuff was about, it is totally unnecessary

OpenStudy (anonymous):

thanks, that makes more sense to me lol

OpenStudy (abb0t):

Idk, I got -0.5 using log. Lol